# FuSeFL: Fully Secure and Scalable Cross-Silo Federated Learning

### v2

@@ V-E The Memory Complexity
Table III summarizes the number of model instances maintained by different entities under AriaNN-FL and FuSeFL (in both Serial and Parallel modes). In AriaNN-FL, the server is involved in both secure training and aggregation, requiring it to maintain $n$ local model instances—one per client—plus the global model. Similarly, the designated aggregator client stores $n$ model shares for aggregation, resulting in significant memory overhead concentrated on central parties. In contrast, FuSeFL distributes the training workload across client pairs, thereby decentralizing memory usage. Each of the two aggregation servers stores only $(n/2)+1$ model instances, cutting the server-side memory requirement by nearly 50%. On the client side, FuSeFL-Serial maintains a single model per group, yielding comparable memory usage to AriaNN-FL but with distributed storage. FuSeFL-Parallel trades memory for speed: each client trains an independent local model copy, increasing client-side memory usage to $3n$ total model instances across all clients. This redundancy enables parallel execution and accelerates training without incurring additional server load.
TABLE III: Number of instances of the model which is stored on different parties. FuSeFL almost halves the number of instances required to be stored on the server.
@@
V-F Security Guarantees
FuSeFL provides strong security guarantees for both data and model confidentiality in a distributed, partially untrusted environment. These guarantees are realized through a combination of secret sharing, secure multi-party computation (MPC), and adversary-aware system design, as defined in our threat model (Section III). Below, we detail the key properties and the cryptographic principles underpinning them.
@@ V-F Security Guarantees
  • Correctness of Secure Training: FuSeFL builds directly on the AriaNN protocol, which guarantees the correctness of MPC-based training among two clients. Each party’s local execution follows the prescribed sequence of arithmetic and Boolean operations as encoded in the secure computation circuit. The resulting model update is functionally equivalent to that of plaintext training. For formal analysis, we refer to the proofs provided in [25].
@@
  • Protection Against Curious Clients: Clients never access the raw data or model parameters of their partner. Instead, they hold one share of their partner’s input and model, making reconstruction infeasible under an honest-but-curious assumption. This ensures that a semi-honest client cannot learn the counterpart’s training data or model parameters. Note that FuSeFL does not assume malicious collusion between group members; however, privacy guarantees degrade proportionally with the number of colluding parties.
@@ V-G Flexibility and Generalization
FuSeFL is designed to be modular with respect to the underlying secure computation protocol. While our current implementation leverages the AriaNN scheme for secure collaborative training within client groups, FuSeFL’s architecture is not bound to this specific backend. The abstraction of secure local training allows FuSeFL to seamlessly generalize to a wider class of MPC schemes.
For instance, adopting a three-party MPC protocol such as SecureNN [23] or Falcon [24] enables FuSeFL to extend its grouping strategy to triads of clients. In such a configuration, each client secret-shares its data and model with the remaining clients in the group, maintaining one share locally. Training proceeds securely over these shared inputs without revealing sensitive data to any single party. The model itself must also be secret-shared in alignment with the MPC protocol’s trust assumptions.
To maintain end-to-end confidentiality, the aggregation infrastructure must scale accordingly. Specifically, FuSeFL requires one aggregation server per client in the group to receive and process the respective secret shares. As a result, for a group of $n$ clients using an $n$-party MPC protocol, FuSeFL can tolerate up to n1n{-}1 colluding clients or servers without compromising data or model confidentiality.
This flexible design allows FuSeFL to support a broad range of MPC backends with different trust and performance trade-offs, making it suitable for deployment in diverse federated learning environments where regulatory, computational, or adversarial constraints may vary.
@@ VI Evaluation Methodology
To assess the performance and accuracy of FuSeFL, we adopt the same experimental setup and neural network architectures used in the AriaNN benchmark [25]. Specifically, we evaluate FuSeFL against AriaNN using three network topologies of varying complexity:
  • Network-1: A simple multilayer perceptron composed of three fully connected layers, each followed by ReLU activation functions.
  • Network-2: A convolutional neural network (CNN) with two convolutional layers (with ReLU and max pooling) followed by two fully connected layers.
  • LeNet: A classic CNN architecture similar to Network-2 but with different layer dimensions and configurations.
All models are trained on the MNIST dataset [48], which is uniformly partitioned among all participating clients. Each client receives an equal share of the data and contributes a single vote to the federated average (i.e., FedAvg weights are uniform across clients). Training configurations and hyperparameters for each network are provided in Table IV.
TABLE IV: Training configurations for different networks
Network # Layers Dataset Learning Rate Momentum
Network1 3 MNIST 0.01 0.9
Network2 4 MNIST 0.01 0.9
LeNet 4 MNIST 0.02 0.9
Experiments are conducted on a machine equipped with an Intel(R) Xeon(R) W-2123 CPU (3.60 GHz, 4 cores). Our FuSeFL implementation is built on top of AriaNN’s implementation [25], which performs secure training using two workers while keeping the data and model private. In our setup, both workers and the coordinating server run as separate processes on the same physical machine, communicating over a local network (LAN). In FuSeFL, the AriaNN training protocol is applied within each client group to collaboratively train the local model.
Figure 8: Accuracy of different federated learning schemes with four local epochs and various global epochs.
@@ VII Evaluation Result
This section evaluate FuSeFL across multiple dimensions, including model accuracy, communication overhead, memory footprint, and training time. We use three baselines for comparison: (1) AriaNN, the original one-server, one-client secure training protocol proposed by Ryffel et al. [25], (2) AriaNN-FL, the extension of AriaNN to a federated learning setting, and (3) WW-FL, a fully offloaded secure federated learning scheme [20].
VII-A Accuracy Evaluation
Table V presents a comparison of test accuracy for FuSeFL and AriaNN when training Network1, Network2, and LeNet on the MNIST dataset, evenly partitioned across 16 clients. AriaNN achieves 98% accuracy on Network1 after 15 epochs of centralized secure training. However, extending AriaNN to a federated setting introduces accuracy degradation: with 4 global and 4 local epochs, AriaNN-FL and FuSeFL-Parallel reach 96.40%, while FuSeFL-Serial achieves 97.20%.
This accuracy gain in FuSeFL-Serial stems from its sequential intra-group training: each group first trains on one client’s data, then continues training on the second client’s data before aggregation. Aggregation is performed per group rather than per client, reducing data fragmentation.
TABLE V: Accuracy of different schemes across different networks
Scheme Epochs* Dataset Network Accuracy(%)
AriaNN[25] 15 MNIST Network1 98.00
AriaNN-FL[25] 4/4 MNIST Network1 96.40
FuSeFL-Parallel 4/4 MNIST Network1 96.40
FuSeFL-Serial 4/4 MNIST Network1 97.20
AriaNN[25] 10 MNIST Network2 98.40
AriaNN-FL[25] 3/4 MNIST Network2 97.80
FuSeFL-Parallel 3/4 MNIST Network2 97.80
FuSeFL-Serial 3/4 MNIST Network2 98.40
AriaNN[25] 10 MNIST LeNet 99.20
AriaNN-FL[25] 4/4 MNIST LeNet 99.00
FuSeFL-Parallel 4/4 MNIST LeNet 99.00
FuSeFL-Serial 4/4 MNIST LeNet 99.20
@@ VII-A Accuracy Evaluation
Importantly, FuSeFL-Serial achieves comparable accuracy $1.43\times$ faster than AriaNN-FL and $6.85\times$ faster than AriaNN. For Network2 and LeNet, FuSeFL-Serial matches AriaNN’s accuracy (98.40% and 99.20%, respectively), demonstrating that FuSeFL delivers the same learning quality as traditional secure training, while offering substantial speedups due to its decentralized design. As observed in Network1, AriaNN-FL and FuSeFL-Parallel yield slightly lower accuracy for Network2 and LeNet, reaching 97.80% and 99.00%, respectively.
@@
VII-B Server-Side Scalability Analysis
To assess FuSeFL’s server-side scalability, we evaluate its ability to handle an increasing number of clients under fixed hardware resources, focusing on the server’s computational overhead for training and aggregation.
As illustrated in Fig. 9, AriaNN-FL’s training time per epoch grows significantly with the number of clients, due to its centralized training coordination. For example, with 128 clients, FuSeFL-Serial achieves up to a $5.65\times$ speedup over AriaNN-FL, while FuSeFL-Parallel achieves up to $7.83\times$. The bottleneck in AriaNN-FL arises from limited parallelism: if a server has 16 compute threads, it must process clients in sequential batches, e.g., training 64 clients in four rounds of 16. This serialization leads to latency that scales linearly with client count, severely limiting scalability.
FuSeFL avoids this bottleneck by shifting training to client pairs. Servers are only responsible for secure aggregation of secret-shared model updates, not training. Since client-group training occurs in parallel and independently of server compute capacity, FuSeFL’s training time remains nearly constant even as the number of clients increases. This design enables scalable, low-latency training suitable for large-scale cross-silo deployments.
Figure 9: Training time per epoch across different schemes as the number of clients increases.
To further highlight the server-side bottleneck, Fig. 10 compares the training throughput, measured as the number of clients trained in parallel, for AriaNN-FL and FuSeFL. In AriaNN-FL, throughput initially increases with client count but quickly saturates once the number of clients exceeds the server’s available CPU cores. This plateau reflects limited server parallelism and growing serialization overhead as more clients participate. In contrast, FuSeFL decouples training from server compute constraints. Since training is offloaded to client pairs, throughput scales linearly with the number of clients, demonstrating FuSeFL’s superior scalability under high client loads.
Figure 10: Training throughput per round. FuSeFL’s throughput is not limited by the server’s computational capabilities during local training
Figure 11: Communication Overhead in FuSeFL vs. AriaNN-FL. By distributing training across client groups, FuSeFL spreads the bandwidth load across the network, substantially reducing client-to-client (C2C) communication overhead.
VII-C Communication Latency at Scale
To assess communication efficiency, we measure the total communication time for one training epoch of Network1 under AriaNN-FL and FuSeFL, assuming each client holds a dataset of 0.45 MB. We break down the overhead into two channels: client-to-client (C2C) and server-to-client/client-to-server (S2C/C2S), as shown in Fig. 11.
In AriaNN-FL (Fig. 11(a)), the server is actively involved in training each client’s model via secure computation, requiring bidirectional communication with all clients per epoch. As the number of clients increases, this centralized setup leads to a linear increase in both S2C/C2S and C2C communication. With 100,000 clients, the overall communication overhead exceeds 11,000 seconds, driven by server-side congestion and sequential communication demands. Even under high parallelism, the server remains a bottleneck due to the required per-client secure coordination.
FuSeFL, shown in Fig. 11(b), mitigates this issue by fully decentralizing the training to client pairs. In this design, C2C communication is localized and parallelized across the network. As a result, this overhead remains nearly constant regardless of scale (e.g.,  0.21 seconds even at 10,000 clients), as communication occurs simultaneously within independent groups. S2C/C2S communication in FuSeFL, which only involves the transmission of final model shares to the servers for secure aggregation, scales linearly but at half the rate of AriaNN-FL. This is because only one model per group is uploaded per round, cutting the volume of communication in half. For instance, at 10,000 clients, FuSeFL’s total communication time is only 45.21 seconds compared to 1,123 seconds in AriaNN-FL, yielding a 25× reduction.
This efficiency gain highlights FuSeFL’s ability to scale securely and efficiently across thousands of clients in cross-silo settings, where communication bandwidth and latency are critical deployment constraints.
VII-D Memory Overhead
Aggregator nodes in federated learning must store multiple instances of local model parameters until all client updates are received for weighted aggregation at the end of each epoch [37]. This requirement leads to considerable memory overhead, especially as the number of clients increases. Figure 12 shows the server-side memory usage for storing model instances. In conventional FedAvg, the server cannot incrementally aggregate updates and must retain all client updates simultaneously, causing memory consumption to scale linearly with the number of participants.
Figure 12: Memory footprint of model instances at the server. FuSeFL reduces server-side memory usage by 50% compared to AriaNN-FL by offloading model training to client groups, thereby minimizing the number of model instances the server must store.
VII-E Comparison with WW-FL[20]
To compare FuSeFL with WW-FL [20], we configure both systems to use AriaNN as their underlying MPC protocol and evaluate them under identical training parameters shown in Table VI. Notably, FuSeFL performs local training without relying on third-party computation servers, whereas WW-FL requires external servers within each cluster to train secret-shared client data.
TABLE VI: Training parameters of $W$-FL and FuSeFL
Parameter $W$-FL FuSeFL
# clients 100 100
# clients per round 100 100
# MPC clusters 10 N/A
# clients per clusters 10 N/A
# clients per clusters per round 10 N/A
Size of clients dataset (MB) 1 1
# Servers used for aggregation 2 2
# Servers used for training 20 0
Note: N/A means FuSeFL does not have the concept of clusters.
Table VII compares the communication cost per training round. In WW-FL, each client must offload both secret shares of its data to designated servers for secure computation, doubling the communication overhead. In FuSeFL, each client retains one share locally and sends the other to its group partner, reducing total communication by 50%. Additionally, the critical communication path, defined as the most bandwidth, congested channel—is substantially lighter in FuSeFL (2 MB vs. 10 MB), improving scalability.
TABLE VII: Comparison of FuSeFL with WW-FL for one round of training
Parameter $W$-FL FuSeFL
Total uploaded data (MB) 200 100
Uploaded data per client (MB) 2 1
Critical communication path (MB) 10 2
WW-FL introduces a trade-off between model accuracy and communication overhead. As the number of MPC clusters decreases, each cluster includes more clients, increasing model accuracy but requiring more data to be transferred per client. Table VIII shows that accuracy improves from 93.00% (C=16) to 96.30% (C=4), but data transfer rises from $N$ to $4N$ per client, undermining FL’s goal of minimizing communication with centralized servers.
In contrast, FuSeFL achieves 95.10% accuracy with only $2N$ data transfer, matching WW-FL with $C=8$ but with no third-party servers required. Moreover, as illustrated in Fig. 8, FuSeFL can achieve further accuracy gains by increasing global epochs without increasing the communication burden.
TABLE VIII: Comparison of FuSeFL with WW-FL in terms of accuracy and communication
$W$-FL FuSeFL
C=16 C=8 C=4
Accuracy 93.00 95.10 96.30 95.10
User Data Transfer N 2N 4N 2N
Unlike WW-FL, which assumes trusted computation servers and requires offloading full data shares for training, FuSeFL retains one share locally and sends only the other to a peer, eliminating the need for a trusted zone. This not only strengthens data privacy but also simplifies deployment in real-world cross-silo settings. Furthermore, by decentralizing secure training to the client layer, FuSeFL avoids WW-FL’s server-heavy design—drastically reducing infrastructure costs, minimizing inter-zone latency, and enabling more scalable, lightweight deployments.
VII-F Robustness Under Threat Model

### v3

@@ V-E The Memory Complexity
Table II summarizes the number of model instances maintained by different entities under AriaNN-FL and FuSeFL (in both Serial and Parallel modes). In AriaNN-FL, the server is involved in both secure training and aggregation, requiring it to maintain $n$ local model instances—one per client—plus the global model. Similarly, the designated aggregator client stores $n$ model shares for aggregation, resulting in significant memory overhead concentrated on central parties. In contrast, FuSeFL distributes the training workload across client pairs, thereby decentralizing memory usage. Each of the two aggregation servers stores only $(n/2)+1$ model instances, cutting the server-side memory requirement by nearly 50%. On the client side, FuSeFL-Serial maintains a single model per group, yielding comparable memory usage to AriaNN-FL but with distributed storage. FuSeFL-Parallel trades memory for speed: each client trains an independent local model copy, increasing client-side memory usage to $3n$ total model instances across all clients. This redundancy enables parallel execution and accelerates training without incurring additional server load.
V-F Security Guarantees
FuSeFL provides strong security guarantees for both data and model confidentiality in a distributed, partially untrusted environment. These guarantees are realized through a combination of secret sharing, MPC, and adversary-aware system design, as defined in our threat model (Section III). Below, we detail the key properties and the cryptographic principles underpinning them.
TABLE II: Number of model instances stored on various parties.
@@ V-F Security Guarantees
  • Correctness of Secure Training: FuSeFL builds directly on the AriaNN protocol, which guarantees the correctness of MPC-based training among two clients. Each party’s local execution follows the prescribed sequence of arithmetic and Boolean operations in the secure computation circuit. The resulting model update is functionally equivalent to that of plaintext training. For formal analysis, we refer to the proofs provided in [22].
@@
  • Protection Against Curious Clients: Clients never access the raw data or model parameters of their partner. Instead, they hold one share of their partner’s input and model, making reconstruction infeasible under an honest-but-curious assumption. This ensures that a semi-honest client cannot learn the counterpart’s training data or model parameters. FuSeFL’s security model assumes that at least one aggregation server is trusted and non-colluding, ensuring that no collusion occurs between the two servers. Inspired by the concept of mix networks [38, 39], this trusted server is additionally responsible for anonymously pairing clients during each training round. By routing all inter-client communication through the trusted server, FuSeFL prevents clients from directly identifying or interacting with their training partners, thereby mitigating the risk of collusion and information leakage.
@@ V-G Flexibility and Generalization
FuSeFL is designed to be modular with respect to the underlying secure computation protocol. While our current implementation leverages the AriaNN scheme for secure collaborative training within client groups, FuSeFL’s architecture is not bound to this specific backend. The abstraction of secure local training allows FuSeFL to seamlessly generalize to a wider class of MPC schemes. For instance, adopting a three-party MPC protocol such as SecureNN [20] or Falcon [21] enables FuSeFL to extend its grouping strategy to triads of clients. In such a configuration, each client secret-shares its data and model with the remaining clients in the group, maintaining one share locally. Training proceeds securely over these shared inputs without revealing sensitive data to any single party. The model itself must also be secret-shared in alignment with the MPC protocol’s trust assumptions. To maintain end-to-end confidentiality, the aggregation infrastructure must scale accordingly. Specifically, FuSeFL requires one aggregation server per client in the group to receive and process the respective secret shares. This flexible design allows FuSeFL to support a broad range of MPC backends with different trust and performance trade-offs, making it suitable for deployment in diverse FL environments where regulatory, computational, or adversarial constraints may vary.
@@ VI Evaluation Methodology
To assess the performance and accuracy of FuSeFL, we adopt the same experimental setup and neural network architectures used in the AriaNN benchmark [22]. Specifically, we evaluate FuSeFL against our baselines using three network topologies:
  • Network-1: A multilayer perceptron composed of three fully connected layers, each followed by ReLU activation functions.
  • Network-2 and LeNet: A convolutional neural network (CNN) with two convolutional layers (with ReLU and max pooling) followed by two fully connected layers.
MNIST dataset [57] is used to train these networks. This dataset is uniformly partitioned among all participating clients. Each client receives an equal share of the data and contributes a single vote to the federated average (i.e., FedAvg weights are uniform across clients). Experiments are conducted on a machine equipped with an Intel(R) Xeon(R) W-2123 CPU (3.60 GHz, 4 cores). Our FuSeFL implementation is built on top of AriaNN’s implementation [22], which performs secure training using two workers while keeping the data and model private. In our setup, both workers and the coordinating server run as separate processes on the same physical machine, communicating over a local network (LAN). In FuSeFL, the AriaNN training protocol is applied within each client group to collaboratively train the local model.
@@ VII Evaluation Result
This section evaluates FuSeFL across multiple dimensions, including model accuracy, communication overhead, memory footprint, and training time. We use three baselines for comparison: (1) AriaNN, the original one-server, one-client secure training protocol proposed by Ryffel et al. [22], (2) AriaNN-FL, the FL extension of AriaNN, and (3) WW-FL, a fully offloaded secure FL scheme [18].
VII-A Server-Side Scalability Analysis
To assess FuSeFL’s server-side scalability, we evaluate its ability to handle an increasing number of clients under fixed hardware resources, focusing on the server’s computational overhead for training and aggregation. As illustrated in Fig. 8(a), AriaNN-FL’s training time per epoch grows significantly with the number of clients, due to its centralized training coordination. For example, with 16 K clients and a dataset of size 46 MB per client, FuSeFL-Serial achieves up to a $13.50\times$ speedup over AriaNN-FL, while FuSeFL-Parallel achieves up to $13.60\times$. The bottleneck in AriaNN-FL arises from limited parallelism: if a server has 16 compute threads, it must process clients in sequential batches, e.g., training 64 clients in four rounds of 16. This serialization leads to latency that scales linearly with client count, severely limiting scalability. FuSeFL avoids this bottleneck by shifting training to client pairs. Servers are only responsible for secure aggregation and routing, not training. Since client-group training occurs in parallel and independently of server compute capacity, FuSeFL’s training time increases steadily as the number of clients increases. This design enables scalable, low-latency training suitable for large-scale cross-silo deployments.
To compare WW-FL [18] with AriaNN and FuSeFL, we configure this system to use AriaNN as its underlying MPC protocol and evaluate them under identical training parameters. Notably, FuSeFL performs local training without relying on third-party computation servers, whereas WW-FL requires external servers within each cluster to train secret-shared client data. As illustrated in Fig. 8(a), WW-FL’s training time increases linearly with the number of clients. This is because more clients are assigned to each server during training. Although the communication overhead is distributed across the network, computation can become the bottleneck as the number of clients grows. For example, with 16 K clients, a dataset size of 46 MB per client, and 16 server clusters, FuSeFL-Serial and FuSeFL-Parallel achieve up to a $9.50\times$ and $9.60\times$ speedup over WW-FL, respectively.
Figure 8: (a) Training throughput per round. FuSeFL’s throughput is not limited by the server’s computational capabilities. (b)Training time per epoch across different schemes as the number of clients increases.
Figure 9: Detailed training time per epoch across different schemes. In FuSeFL, computation time is negligible when increasing the number of clients.
As shown in Fig. 9, FuSeFL’s end-to-end training time becomes increasingly dominated by communication latency as the number of clients increases, while computation latency remains comparatively small. This is because computation is distributed among clients rather than a limited number of servers. In contrast, AriaNN-FL and WW-FL incur substantially higher total training times; although their communication overhead is lower than FuSeFL’s, their computation overhead becomes the bottleneck at larger client counts. Adding a trusted server as a routing hop between the clients results in an increase in the communication time in FuSeFL, which is approximately $2\times$ and $32\times$ higher than that of AriaNN and WW-FL, respectively. However, by reducing the computation time, FuSeFL can achieve a better overall training time.
To further highlight the server-side bottleneck, Fig. 8(b) compares the training throughput, measured as the number of clients trained in parallel, for AriaNN-FL, WW-FL, and FuSeFL. In both AriaNN-FL and WW-FL, throughput initially increases with client count but quickly saturates once the number of clients exceeds the server’s available CPU cores (in AriaNN-FL) or the number of server clusters (in WW-FL). This plateau reflects limited server parallelism and growing serialization overhead as more clients participate. In contrast, FuSeFL decouples training from server compute constraints. Since training is offloaded to client pairs, throughput scales linearly with the number of clients, demonstrating FuSeFL’s superior scalability under high client loads.
Figure 10: Communication Overhead in FuSeFL vs. AriaNN-FL.
VII-B Communication Latency at Scale
We measure the total communication time for one training epoch of Network1 under AriaNN-FL and FuSeFL, assuming each client holds a dataset of 46 MB. We break down the overhead into two channels described on §V-D: C2C and S2C/C2S, as shown in Fig. 10.
FuSeFL’s communication overhead exceeds that of AriaNN-FL due to the need to support non-colluding parties. Inspired by the idea of mixnets, we propose using a trusted node (one of the servers in our case) as a bridge to connect two computing parties anonymously. As shown in Fig. 10(b), in this design, C2C communication is routed through the trusted server. As a result, this overhead increases linearly with the number of clients. S2C/C2S communication in FuSeFL, which only involves transmitting the final model shares to the servers for secure aggregation, also scales linearly but at half the rate of AriaNN-FL. This is because only one model per group is uploaded per round, cutting the communication volume in half. It is true that FuSeFL’s communication latency is twice that of AriaNN-FL, but the benefits FuSeFL gains by distributing the training cancel out this drawback, specifically, without losing any security guarantees. This efficiency gain highlights FuSeFL’s ability to scale securely and efficiently across thousands of clients. Although FuSeFL employs a mixnet-style trusted routing layer to anonymize communication between clients, the additional delay introduced by this mechanism is negligible. For example, in our experiments with 256 clients, the per-pair routing and shuffling delay was approximately 2% of the total communication time and did not affect overall training throughput. Thus, the mixnet integration in FuSeFL provides robust privacy guarantees with minimal performance cost.
VII-C Memory Overhead
Aggregator nodes in federated learning must store multiple instances of local model parameters until all client updates are received for weighted aggregation at the end of each epoch [37]. This requirement leads to considerable memory overhead, especially as the number of clients increases. Figure 11 shows the server-side memory usage for storing model instances for Network1. In conventional FedAvg, the server cannot incrementally aggregate updates and must retain all client updates simultaneously, causing memory consumption to scale linearly with the number of participants. In WW-FL, however, the memory usage on the server side remains steady as the number of clients increases. This amount is limited by the number of server clusters used for training. On the server, we need to store one instance of the model per server cluster.
Figure 11: Memory footprint of model instances at the server. FuSeFL reduces server-side memory usage by 50% compared to AriaNN-FL by offloading training to client groups.
Figure 12: Accuracy of different FL schemes for various global epochs and 4 local epochs.
TABLE III: Accuracy comparison of different schemes
Scheme Epochs* Network Accuracy(%)
AriaNN[22] 15 Network1 98.00
AriaNN-FL[22] 4/4 Network1 96.40
FuSeFL-Parallel 4/4 Network1 96.40
FuSeFL-Serial 4/4 Network1 97.20
AriaNN[22] 10 Network2 98.40
AriaNN-FL[22] 3/4 Network2 97.80
FuSeFL-Parallel 3/4 Network2 97.80
FuSeFL-Serial 3/4 Network2 98.40
AriaNN[22] 10 LeNet 99.20
AriaNN-FL[22] 4/4 LeNet 99.00
FuSeFL-Parallel 4/4 LeNet 99.00
FuSeFL-Serial 4/4 LeNet 99.20
@@ VII-C Memory Overhead
VII-D Accuracy Evaluation
Table III presents a comparison of test accuracy for FuSeFL and AriaNN when training Network1, Network2, and LeNet on the MNIST dataset, evenly partitioned across 16 clients. AriaNN achieves 98% accuracy on Network1 after 15 epochs of centralized secure training. However, extending AriaNN to a federated setting introduces accuracy degradation: with 4 global and 4 local epochs, AriaNN-FL and FuSeFL-Parallel reach 96.40%, while FuSeFL-Serial achieves 97.20%. This accuracy gain in FuSeFL-Serial stems from its sequential intra-group training: each group first trains on one client’s data, then continues training on the second client’s data before aggregation. Aggregation is performed per group rather than per client, reducing data fragmentation. Importantly, FuSeFL-Serial achieves comparable accuracy $1.43\times$ faster than AriaNN-FL and $6.85\times$ faster than AriaNN. For Network2 and LeNet, FuSeFL-Serial matches AriaNN’s accuracy (98.40% and 99.20%, respectively), demonstrating that FuSeFL delivers the same learning quality, while offering substantial speedups due to its decentralized design. AriaNN-FL and FuSeFL-Parallel yield slightly lower accuracy for Network2 and LeNet, reaching 97.80% and 99.00%, respectively.
@@ VII-D Accuracy Evaluation
VII-E Robustness Under Threat Model